A 12 N force with a fixed orientation does work on
Answer #1

given

F = 12 N

d = 4 i - 5 j + 3 k

+ 8.49J , - 8.49 J

a )

cos gif.latex?%5Csmall%20%5Ctheta = (+8.49) / 12 ( 42 + (-5)2 + 32 )

cos gif.latex?%5Csmall%20%5Ctheta = 8.49 / 12 ( 50 )

cos gif.latex?%5Csmall%20%5Ctheta = 8.49 / 600

cos gif.latex?%5Csmall%20%5Ctheta = 0.01415

gif.latex?%5Csmall%20%5Ctheta = cos-1(0.01415)

\small \theta= 89.1890

b )

cos gif.latex?%5Csmall%20%5Ctheta = (-8.49) / 12 ( 42 + (-5)2 + 32 )

cos gif.latex?%5Csmall%20%5Ctheta = -8.49 / 12 ( 50 )

cos gif.latex?%5Csmall%20%5Ctheta = - 8.49 / 600

cos gif.latex?%5Csmall%20%5Ctheta = - 0.01415

gif.latex?%5Csmall%20%5Ctheta = cos-1(-0.01415)

\small \theta= 90.810